# Supervised Fine-Tuning: An Activation Pattern Optimization Process for Attention Heads

### v1

@@ Results
As shown in Table 1, before SFT, the base models exhibited significant outliers (Kurtosis ¿ 3), uneven distribution (Gini ¿ 0.4), and high variability (CV ¿ 1) in activation patterns across multiple tasks. This suggests that only a few attention heads have high activation levels and are contributing to the task. After the SFT process, the distribution of activation patterns became more uniform, with decreases observed in the Gini, CV, and Kurtosis index. This suggests that SFT adapts LLMs to specific tasks by increasing the activate level of attention heads. However, the distribution of LLMs activation values remains highly skewed, reflecting that only a few heads play a significant role.
@@
Figure 2 shows the similarity in the activation pattern change rates across different tasks for three models before and after SFT, measured by correlation coefficients. Moreover, as shown in Figure 2, based on the correlation analysis, after SFT, the tasks were divided into two categories: math and code-related tasks, and text reasoning tasks. This largely aligns with human understanding of the relationships between these tasks. Therefore, the attention patterns can, to some extent, reflect the characteristics of the tasks and exhibit task specificity. After SFT, the changes in the model’s activation patterns demonstrate stronger task specificity, as evidenced by the decrease in the correlation coefficients of the attention patterns across different tasks—in other words, the similarity between tasks has decreased.
@@
Activation Patterns on Complex Tasks are Combinations of Basic Tasks
@@
In this paper, we define complex tasks as those that can be decomposed into more basic skills. For instance, as illustrated in Figure 3, solving the ”chicken and rabbit” problem requires elementary math skills, while constructing a binary tree requires Python coding skills—both considered basic tasks. However, solving the ”chicken and rabbit” problem using Python code is a complex task, as it requires both mathematical reasoning and coding skills.
@@
Experiments Setup
In this study, we conducted two groups of experiments: (1) We take MATH, GSM8K, Code Search Net, HellaSwag, Winogrande, and ARC as the basic tasks, and select the SGSM task, which requires coding ability and elementary math skills, as the complex task. (2) We conduct experiments on the Infinity Instruct dataset (of Artificial Intelligence BAAI), which provides each instruction tags describing the skills and knowledge required to complete that instruction, so that we can select both basic and complex task-related instructions. Specifically, we selected instructions requiring only a single skill, including ‘mathematics’, ‘logical reasoning’, ‘programming and software development’, ‘education and consulting’, ‘financial and business knowledge’, and ‘legal knowledge’, to form into the basic task data set. Next, we selected instructions requiring both ‘logical reasoning’ and ‘programming/software development’ skills as the complex tasks for SFT and calculated their activation patterns.
Result
@@
Small Data Can Change Activation Patterns
@@
As shown in Figure 5, activation patterns change rapidly in the initial steps of SFT. This suggests that fine-tuning with small datasets can significantly reshape a model’s attention patterns and alter its performance. Previous research (Xia et al. 2024) supports this, indicating that rapid learning during SFT is driven by swift changes in attention patterns.
@@
Scenario 1: Combining Data of Basic Tasks to Promote the Learning of Complex Task
@@
Given that the relationships between the complex task and basic tasks are unknown, we first fit these relationships using the regression function described in Equation 3. Intuitively, the regression coefficients $\alpha_{i}$ describe the proportion that the complex task can be “composed ” by a basic task $i$. Thus, we construct a preliminary instruction set by mixing instructions about basic tasks according to the magnitude of the regression coefficients:
$\text{Dataset}^{preliminary}=\{N\times\alpha_{i}/\sum_{i}\alpha_{i}\}_{i=0}^{|B|}$ (4)
where $|B|$ is the number of candidate basic tasks, $N$ is the predefined size of $\text{Dataset}^{preliminary}$. During experiment, LLM is first trained on $\text{Dataset}^{preliminary}$, then finetuned on the target task dataset $\text{Dataset}^{\text{complex task}}$.
Experiment Settings
To simulate scenarios where LLMs lack sufficient preliminary knowledge, we conducted experiments on the Llama series models (Touvron et al. 2023a, b; AI@Meta 2024), which have not been specifically optimized for Chinese corpora and exhibit relatively limited performance on Chinese-related tasks. In this section, we set the target complex task as MathBench (Liu et al. 2024), which requires the model to solve school mathematics questions using Chinese. The MathBench dataset 358 instances, and we use only 100 instances as $\text{Dataset}^{\text{complex task}}$, the left part for testing. In addition to the basic tasks described in Section 4.2, we include RefGPT (Yang et al. 2023) (a dataset of Chinese factual knowledge) as a basic task dataset. After the regression process, we keep only basic datasets with the Top 2 largest regression coefficients, so as to filter out the noises. The remaining basic tasks are GSM8K (English elementary math) and RefGPT.Experiments are conducted on models of similar size but varying capabilities: Llama-7B (Touvron et al. 2023a), Llama2-7B (Touvron et al. 2023b), and Llama3-8B (AI@Meta 2024). Detailed experimental procedures are in Appendix I and II.
@@
Table 2: Performance on the Mathbench test set: ”Base” refers to the base model results, “SFT” refers to results after fine-tuning with 100 Mathbench samples, and “Random” represents the average of five runs trained with a randomly mixed dataset of GSM8K and RefGPT data, followed by fine-tuning with 100 Mathbench samples Best results are bolded, and second-best are underlined.
Analysis
As shown in Table 2, our approach consistently outperforms baseline models across all evaluations, including models fine-tuned solely on the Mathbench training set and those trained with a random mix of elementary mathematics and Chinese instruction data. Compared to models fine-tuned only on Mathbench, our method shows significant improvements in handling complex tasks through subtask training. Additionally, using activation pattern proportions for dataset construction enhances the model’s ability to utilize simple task data for complex tasks, outperforming models trained on randomly combined subtasks. For Llama-7B, learning from randomly proportioned basic task data can still improve complex task performance, but for more powerful models like Llama2-7B and Llama3-8B, this approach may hinder performance. These findings suggest that guiding dataset proportioning with activation patterns from basic tasks significantly enhances the model’s ability to learn and generalize to complex tasks, reinforcing our earlier analysis of the linear relationship between basic and complex tasks.
Scenario 2: Selecting Relevant Data from Candidates
While various open-source Domain LLMs demonstrate strong capabilities, their instruction data is often unavailable. Hence, considering the task-specificity of activation patterns, could we select relevant datasets from open instruction sets and use them to approximate the performance trained upon private domain datasets? Thus, we conduct experiments under the following conditions: (1) only a small amount of developer-constructed pseudo-private data is available;(2) the complete dataset used to train the domain model is unknown;(3) The public data is sufficiently large, and there exists data that can improve the model’s performance on the target task.
Experiment Settings
@@
Table 3: Average accuracy of the model on four tasks: mathematics, physics, chemistry, and biology. ”Base” represents the baseline model’s average accuracy on these tasks. “T100” indicates the accuracy after fine-tuning with 100 most similar data points based on activation patterns. ”R100” represents the accuracy after fine-tuning with 100 randomly selected unrelated MMLU data points, with similar meanings for “T300”, “R300”, “T500”,and “R500”.
To simulate the performance of LLMs when target data is unavailable, we conducted experiments on the MMLU dataset (Hendrycks et al. 2021a), which covers multiple domains. We selected test set data from certain domains as private (target data) and treated the remaining MMLU test set data as non-target data. Additionally, we used validation and dev set data from the same domains as pseudo-private data, simulating the small, manually constructed datasets that developers might create when the original data is unavailable. Fine-tuning on this pseudo-private data enhanced the model’s task specificity for the target domain. We then used Equation 5 to identify the top $m$ non-target data points from MMLU that are most similar to the pseudo-private data, and used these for further training. We compared the performance of this approach with the base model and models trained on an equal amount of randomly selected non-target data. To ensure robustness, we averaged the results across multiple domains, including mathematics, physics, chemistry, and biology, treating them as target data.
$\text{Dataset}^{approx}=\operatorname{Top}_{m}Corr(AP^{target},AP^{D_{i}}),$ (5)
@@
where $\text{Dataset}^{approx}$ represents the entire dataset, $D_{i}$ is the $i$-th sample from $D$, $AP^{D_{i}}$ is the activation pattern for the $i$-th sample after SFT, and $Corr(*,*)$ is the correlation coefficient.
Analysis
@@
As shown in Table 3, models trained with data selected by activation pattern similarity consistently outperform those trained with randomly selected data. This suggests that using the similarity of activation patterns, we can effectively identify data that share similarities with the private data.These experimental results not only support our theoretical hypothesis regarding activation patterns but also demonstrate the broad potential of this strategy in practical applications. By judiciously selecting training data based on activation patterns, we can achieve significant performance improvements in scenarios where target data is unavailable, providing a robust methodological approach for tackling complex tasks.
Conclusion
@@
This study investigated the mechanisms behind the rapid learning and generalization observed during SFT. We found that attention heads are selectively activated in a task-specific manner during SFT, and that activation patterns for complex tasks can be decomposed into combinations of basic task patterns, offering a viable data substitution strategy when high-quality complex data is scarce.
@@
Our experiments demonstrated that even small amounts of data can significantly alter activation patterns, validating the feasibility of (1) pre-training on simple tasks to equip large models with essential knowledge, and (2) fine-tuning with selected relevant data to enable rapid generalization. The improved effectiveness observed in these scenarios supports the validity of our analytical approach and offers practical solutions to data challenges in complex and specialized tasks.

### v2

@@ 4.1.3 Results
As shown in Table 1, before SFT, base models exhibited significant outliers (Kurtosis > 3), uneven distribution (Gini > 0.4), and high variability (CV > 1) in activation patterns, suggesting that only a few attention heads contributed to the tasks. After SFT, activation patterns became more uniform, with decreases in Gini, CV, and Kurtosis, indicating that SFT adapts LLMs by increasing activation levels across attention heads. However, activation values remain skewed, showing that a few heads still dominate.
@@
Figure 2 shows the similarity in activation pattern change rates across tasks for three models before and after SFT, measured by correlation coefficients. After SFT, tasks grouped into two categories—math/code and text reasoning—aligning with human understanding of task relationships. This suggests that attention patterns reflect task characteristics and specificity. Post-SFT, stronger task specificity is observed, as indicated by decreased correlation coefficients between tasks, implying reduced similarity.
@@
4.2 Activation Patterns on Complex Tasks are Combinations of Basic Tasks
@@ 4.2.1 Analytical Method
In this paper, we define complex tasks as those that can be decomposed into basic skills. For example, as shown in Figure 3, solving the “chicken and rabbit” problem requires mathematical reasoning, while constructing a binary tree requires Python coding—both of which are basic tasks. However, solving the “chicken and rabbit” problem using Python code is a complex task, as it integrates both mathematical reasoning and coding skills.
@@
4.2.2 Experiments Setup
In this study, we conducted two groups of experiments: (1) We take MATH, GSM8K, Code Search Net, HellaSwag, Winogrande, and ARC as the basic tasks, and select the SGSM task, which requires coding ability and elementary math skills, as the complex task. (2) We conduct experiments on the Infinity Instruct dataset Zhao et al. (2024), which provides each instruction tags describing the skills and knowledge required to complete that instruction, so that we can select both basic and complex task-related instructions. Specifically, we selected instructions requiring only a single skill, including ‘mathematics’, ‘logical reasoning’, ‘programming and software development’, ‘education and consulting’, ‘financial and business knowledge’, and ‘legal knowledge’, to form into the basic task data set. Next, we selected instructions requiring both ‘logical reasoning’ and ‘programming/software development’ skills as the complex tasks for SFT and calculated their activation patterns.
4.2.3 Result
@@
4.3 Small Data Can Change Activation Patterns
@@ 4.3.2 Experiments and Results
As shown in Figure 5, activation patterns change rapidly in the initial steps of SFT. This suggests that fine-tuning with small datasets can significantly reshape a model’s attention patterns and alter its performance. Previous research  Xia et al. (2024) supports this, indicating that rapid learning during SFT is driven by swift changes in attention patterns.
@@ 5 Deduction and Applications
5.1 Scenario 1: Combining Data of Basic Tasks to Promote the Learning of Complex Task
@@
Since the relationships between the complex task and basic tasks are unknown, we first estimate these relationships using the regression function described in Equation 3. Intuitively, the regression coefficients $\alpha_{i}$ describe the proportion that the complex task can be “composed ” by a basic task $i$. Based on the regression coefficients, we create a preliminary instruction set by combining instructions from basic tasks.
$\text{Dataset}^{pre}=\{N\times\alpha_{i}/\sum_{i}\alpha_{i}\}_{i=0}^{|B|}$ (4)
where $|B|$ is the number of candidate basic tasks, $N$ is the predefined size of $\text{Dataset}^{pre}$. During experiment, LLM is first trained on $\text{Dataset}^{pre}$, then finetuned on the target task dataset $\text{Dataset}^{\text{complex}}$.
5.1.1 Experiment Settings
To simulate scenarios where LLMs lack sufficient preliminary knowledge, we conducted experiments on the Llama series models, which have not been specifically optimized for Chinese corpora and exhibit relatively limited performance on Chinese-related tasks. In this section, we set the target complex task as MathBench  Liu et al. (2024), which requires the model to solve school mathematics questions using Chinese. The MathBench dataset 358 instances, and we use only 100 instances as $\text{Dataset}^{\text{complex}}$, the left part for testing. In addition to the basic tasks described in Section 4.2, we include RefGPT  Yang et al. (2023a) (a dataset of Chinese factual knowledge) as a basic task dataset. After the regression process, we keep only basic datasets with the Top 2 largest regression coefficients, so as to filter out the noises. The remaining basic tasks are GSM8K (English elementary math) and RefGPT.Experiments are conducted on models of similar size but varying capabilities: Llama-7B  Touvron et al. (2023a), Llama2-7B  Touvron et al. (2023b), and Llama3-8B.
@@ 5.1.1 Experiment Settings
Table 2: Performance on the Mathbench test set: "Base" refers to the base model results, “SFT” refers to results after fine-tuning with 100 Mathbench samples, and “Random” represents the average of five runs trained with a randomly mixed dataset of GSM8K and RefGPT data, followed by fine-tuning with 100 Mathbench samples.
5.1.2 Analysis
As shown in Table 2, our approach consistently outperforms baseline models across all evaluations, including models fine-tuned solely on the Mathbench training set and those trained with a random mix of elementary mathematics and Chinese instruction data. Compared to models fine-tuned only on Mathbench, our method shows significant improvements in handling complex tasks through subtask training. Moreover, constructing datasets based on activation patterns improves the model’s ability to leverage simple task data for complex tasks, surpassing models trained on randomly mixed subtasks. For Llama-7B, learning from randomly proportioned basic task data can still improve complex task performance, but for more powerful models like Llama2-7B and Llama3-8B, this approach may hinder performance. These findings suggest that guiding dataset construction using activation patterns from basic tasks significantly enhances the model’s ability to learn and generalize to complex tasks, supporting our observation that the complex tasks could be the combination of basic complex tasks, and once equipping necessary preliminary knowledge the LLM could fast adapt to complex task.
5.2 Scenario 2: Selecting Relevant Data from Candidates
While various Domain-LLMs demonstrate strong capabilities, their instruction data is often unavailable. Considering the task-specificity of activation patterns, can we select relevant datasets from open instruction sets to approximate the performance of models trained on private domain datasets? Thus, we conduct experiments under the following settings: (1) only a small amount of developer-constructed pseudo-private data is available; (2) the complete dataset used to train the Domain-LLM is unknown;(3) the public data pool is large enough to contain data relevant to the target task.
5.2.1 Experiment Settings
@@
Table 3: Average accuracy of the model on four tasks: mathematics, physics, chemistry, and biology. "Base" represents the baseline model’s average accuracy on these tasks. “T100” indicates the accuracy after fine-tuning with 100 most similar data points based on activation patterns. "R100" represents the accuracy after fine-tuning with 100 randomly selected unrelated MMLU data points, with similar meanings for “T300”, “R300”, “T500”, and “R500”.
We simulated LLM performance without target data using the multi-domain MMLU dataset, designating test set data from specific domains as private (target data) and treating the remaining as non-target data. We used validation and dev set data from the target domains as pseudo-private data, simulating small, manually created datasets that developers might construct when original data is unavailable. As concluded in Section 4.1, fine-tuning on pseudo-private data improves the model’s task specificity in the target domain.
We used Equation 5 to identify the top $m$ non-target data points from MMLU most similar to the pseudo-private data. These data points were then used for further training. We compared the performance of this approach with the base model and models trained on an equal amount of randomly selected non-target data.
To ensure robustness, we averaged the results across multiple domains—mathematics, physics, chemistry, and biology—treating them as target data.
$\text{Dataset}^{app}=\operatorname{Top}_{m}Corr(AP^{tar},AP^{D_{i}}),$ (5)
@@
where $\text{Dataset}^{app}$ represents the entire dataset, $D_{i}$ is the $i$-th sample from $D$, $AP^{D_{i}}$ is the activation pattern for the $i$-th sample after SFT, and $Corr(*,*)$ is the correlation coefficient.
5.2.2 Analysis
@@
As shown in Table 3, models trained with data selected by activation pattern similarity consistently outperform those trained with randomly selected data. This suggests that using the similarity of activation patterns, we can effectively identify data that share similarities with the private data. These results support our hypothesis that attention heads carry task-specific information, enabling LLMs to adapt to various tasks. They also demonstrate the potential for identifying rare relevant data from large datasets to approximate private datasets.
6 Conclusion
@@
This study investigated the mechanisms behind the rapid learning and generalization observed during SFT. We found that attention heads are selectively activated in a task-specific manner during SFT, and that the activation patterns for complex tasks can be decomposed into combinations of basic task patterns. This offers a viable data substitution strategy when high-quality complex data is scarce.
@@
Our experiments demonstrated that even small amounts of data can significantly alter activation patterns, validating the feasibility of (1) pre-training on simple tasks to equip large models with essential knowledge, and (2) fine-tuning with selected relevant data to enable rapid generalization. The improved effectiveness observed in these scenarios further supports the validity of our analytical approach and offers practical solutions to data challenges in complex and specialized tasks.
7 Limitations
While this study demonstrates the role of attention head activation patterns in rapid learning and generalization during the SFT process, we did not explore the specific impact of individual attention head activation levels on model performance in detail. Additionally, our method validation was primarily conducted on simpler tasks, and its applicability to more complex and real-world tasks remains to be fully evaluated. Future work will focus on finer-grained analysis of activation levels and further testing the effectiveness of our approach in more challenging and practical domains to enhance its performance in real-world applications.
Appendix A Dataset Details
In this paper, we utilize a wide range of datasets related to LLMs, many of which are commonly used in prominent large model evaluation frameworks Fourrier et al. (2024); Contributors (2023).
  • MATH Hendrycks et al. (2021b): An advanced mathematical reasoning dataset.
  • GSM8K Cobbe et al. (2021): A dataset for elementary mathematical reasoning.
  • Code Search Net Husain et al. (2019): A large dataset of code sourced from GitHub; we specifically use the Python subset in this paper.
  • SGSM Christ et al. (2024): A dataset for solving math problems using code, requiring strong code generation and mathematical reasoning abilities.
  • HellaSwag Zellers et al. (2019): A dataset for natural language processing and reasoning tasks.
  • Winogrande Sakaguchi et al. (2021): Another dataset for natural language processing and reasoning tasks.
  • ARC Clark et al. (2018): A dataset for natural language processing and reasoning tasks.
  • Infinity Instruct dataset Zhao et al. (2024): A bilingual dialogue dataset; we use the English portion in this study. This dataset is annotated with the specific skills required to complete each dialogue, allowing us to determine which capabilities are needed for each task.
  • MathBench Liu et al. (2024): A dataset for evaluating mathematical performance in large models. We use the Chinese portions of the middle and primary subsets.
  • RefGPT Yang et al. (2023a): A Chinese dataset focused on factual knowledge.
  • MMLU Hendrycks et al. (2021a): A dataset containing data from 57 domains. Following the approach of Wang et al. (2024b), we use "college mathematics," "abstract algebra," "elementary mathematics," "high school statistics," and "high school mathematics" to represent math-related data; "college physics," "high school physics," and "conceptual physics" for physics-related data; "college chemistry" and "high school chemistry" for chemistry-related data; and "college biology" and "high school biology" for biology-related data.
Appendix B Training Details
  • All experiments were conducted on an NVIDIA A100 GPU cluster.
  • The training was performed using the DeepSpeed framework, following the hyperparameters outlined in Yuan et al. (2023). We set ZeRO to 2, used a batch size of 4, a maximum token length of 1024, a learning rate of 1e-6, and gradient accumulation set to 4. Full fine-tuning was conducted with bf16 precision.
  • Given the varying sizes of the datasets, to ensure consistency within each experiment, the number of training samples was set to match the smallest dataset in the group, ensuring that the results were not influenced by the amount of data.
  • To ensure fair evaluation, all assessments were conducted using the OpenCompass framework Contributors (2023), employing greedy search to eliminate randomness in the results.
Appendix C Equation Details
MSE represents the average distance between two matrices. The larger the MSE, the greater the difference between the two matrices.
$\begin{split}\text{MSE}(\mathbf{AP}^{\text{task1}},\mathbf{AP}^{\text{task2}})&=\frac{1}{L\times H}\sum_{l=1}^{L}\sum_{h=1}^{H}\\ &\quad\left(\mathbf{AP}^{\text{task1}}_{l,h}-\mathbf{AP}^{\text{task2}}_{l,h}\right)^{2},\end{split}$ (6)
where $\mathbf{AP}^{\text{task1}}$ and $\mathbf{AP}^{\text{task2}}$ represent the activation pattern matrices for tasks 1 and 2, respectively, and $L$ and $H$ represent the number of layers and the number of heads per layer, respectively.
The Gini coefficient, which measures the inequality within the data of a matrix $\mathbf{AP}$, can be expressed as:
$G(\mathbf{AP})=\frac{\sum_{i=1}^{L}\sum_{j=1}^{H}\sum_{k=1}^{L}\sum_{l=1}^{H}|\mathbf{AP}_{ij}-\mathbf{AP}_{kl}|}{2LH\sum_{i=1}^{L}\sum_{j=1}^{H}\mathbf{AP}_{ij}}$ (7)
Coefficient of Variation The Coefficient of Variation (CV), which measures the relative variability of the data in matrix $\mathbf{AP}$, is given by:
$CV(\mathbf{AP})=\frac{\sigma(\mathbf{AP})}{\mu(\mathbf{AP})},$ (8)
where $\sigma(\mathbf{AP})$ is the standard deviation of $\mathbf{AP}$, and $\mu(\mathbf{AP})$ is the mean of $\mathbf{AP}$.
The Kurtosis, which measures the peakedness of the data distribution in matrix $\mathbf{AP}$, can be expressed as:
$K(\mathbf{AP})=\frac{LH\sum_{i=1}^{L}\sum_{j=1}^{H}(\mathbf{AP}_{ij}-\mu(\mathbf{AP}))^{4}}{\left(\sum_{i=1}^{L}\sum_{j=1}^{H}(\mathbf{AP}_{ij}-\mu(\mathbf{AP}))^{2}\right)^{2}},$ (9)
where $\mu(\mathbf{AP})$ is the mean of the matrix $\mathbf{AP}$.
These expressions can be used in a LaTeX document to calculate the Gini coefficient, Coefficient of Variation, and Kurtosis for the matrix $\mathbf{AP}$.
We use $R^{2}$ to analyze the task fitting performance:
$R^{2}=1-\frac{\sum_{l=1}^{L}\sum_{h=1}^{H}\left(\Delta AP^{\text{complex}}_{lh}-\Delta\hat{AP}^{\text{complex}}_{lh}\right)^{2}}{\sum_{l=1}^{L}\sum_{h=1}^{H}\left(\Delta AP^{\text{complex}}_{lh}-\overline{\Delta AP^{\text{complex}}}\right)^{2}}$ (10)
The coefficient of determination, denoted as $R^{2}$, is a statistical measure that indicates how well the data fits a regression model. In this context, $\Delta AP^{\text{complex}}_{lh}$ represents the observed change in activation patterns at layer $l$ and head $h$ for a complex task, while $\Delta\hat{AP}^{\text{complex}}_{lh}$ denotes the corresponding predicted change. The numerator captures the sum of squared errors between the observed and predicted values, whereas the denominator reflects the total variance in the observed data. $\overline{\Delta AP^{\text{complex}}}$ represents the mean of the observed changes in activation patterns.